p(x)=x^4+kx^2+9
what are all the possible values of k? what are the real roots?

Answer by stanbon(75887)

p(x)=x^4+kx^2+9
what are all the possible values of k? what are the real roots
---------------------
Solve::
x^4 + kx^2 + 9 = 0
------
If k = -6, (x^2-3)^2 = 0
------
Then x^2 = 3
And x = 3 or x = -3
============================
Cheers,
Stan H.